1.   Molarity (M)

Calculate the molar concentration of a solution containing 13.4 g potassium hydroxide KOH in 355 mL of solution.


1
Expert's answer
2022-03-04T12:04:03-0500

Molar mass of KOH = 56.1056 g/mol

13.4/56.1056= 0.24

= 0.24/0.355= 0.676